Ramesh had 20 pencils, Sheelu had 50 pencils and Jamaal had 80 pencils. After 4 months, Ramesh used up 10 pencils, Sheelu used up 25 pencils and Jamaal used up 40 pencils. What fraction did each use up? Check if each has used up an equal fraction of her/his pencils?

Grade:12

2 Answers

Harshit Singh
2 years ago
Dear Student

Total number of pencils Ramesh had = 20 Number of pencils used by Ramesh = 10
∴Fraction = 10 / 20 = 1 / 2
Total number of pencils Sheelu had = 50 Number of pencils used by Sheelu = 25
∴Fraction = 25 / 50 = 1 / 2
Total number of pencils Jamaal had = 80 Number of pencils used by Jamaal = 40
∴Fraction = 40 / 80 = 1 / 2
Yes, each has used up an equal fraction of pencils i.e 1 / 2

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
No. Of pencils Ramesh have = 20
He used = 10
Fraction used = 10/20 = 1/2
 
No. Of pencils Sheelu have = 50
She used = 25
Fraction used = 25/50 = ½
 
No. Of pencils Jamaal have = 80
He used = 40
Fraction used= 1/2
 
Therefore, each of them have used equal fraction of pencils.
 
Hope it helps!
Thank you ✌️